OpenStudy (anonymous):

Solve the problem algebraically: 4/y - 5 = 18/2y Can someone show me how to work this??

OpenStudy (anonymous):

\[{ 4 \over { y-5} } = { { 18 } \over { 2y} } ?\]

OpenStudy (anonymous):

- 5 is beside 4 over y

OpenStudy (anonymous):

\[{ 4 \over y } -5 = { { 18 } \over { 2y} } ?\] and you need all the steps i presume?

OpenStudy (anonymous):

please and thank you

OpenStudy (pdoctor):

First you would create a common denominator for the left hand side. \[4/y -5y/y = 9/y\] Then combine the left hand side \[(4-5y)/y\] Then Multiply both sides by y \[4-5y=9\] Then subtract 4 \[-5y=5\] Then divide by -5 \[y=-1\]

OpenStudy (anonymous):

thank you very much